A factory manufacturing light-emitting diode (LED) bulbs claims that their light bulb lastfor 50 000 hours on the average. To confirm its this claim is valid, a quality control manager got a sample of 50 LED bulbs and obtained a life span of 40 000 hours. The standard deviation of the manufacturing process is 1 000 hours. Do you think the claim of the manufacturer is valid at the 5% level of significance?

1
Expert's answer
2022-02-15T14:23:49-0500

"\bar{x}=40000 \\

s=1000 \\

n=50 \\

df=n-1=49 \\

α=0.05 \\"

The hypotheses tested are,

"H_0: \mu=50000 \\vs\\

H_1: \mu ≠50000 \\"

The critical value is,

"t_{{\alpha\over 2},df}=t_{0.025,49}= 2.009575"

The rejection region for this two-tailed test is R={t:|t|>2.009575}

Test-statistic:

"t = \frac{\bar{x}- \mu}{s\over \sqrt{n}} \\

t = \frac{40000-50000}{1000 \over \sqrt{50}} = -70.711 \\

|t|=70.711>2.009575"

The null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean is different than 50000, at the 0.05 significance level.